### v1

# Efficient Crystal Structure Prediction Using Genetic Algorithm and Universal Neural Network Potential

Abstract
Crystal structure prediction (CSP) is crucial for identifying stable crystal structures in given systems and is a prerequisite for computational atomistic simulations. Recent advances in neural network potentials (NNPs) have reduced the computational cost of CSP. However, searching for stable crystal structures across the entire composition space in multicomponent systems remains a significant challenge. Here, we propose a novel genetic algorithm (GA) -based CSP method using a universal NNP. Our GA-based methods are designed to efficiently expand convex hull volumes while preserving the diversity of crystal structures. This approach draws inspiration from the similarity between convex hull updates and Pareto front evolution in multi-objective optimization. Our evaluation shows that the present method outperforms the symmetry-aware random structure generation, achieving a larger convex hull with fewer trials. We demonstrated that our approach, combined with the developed universal NNP (PFP), can accurately reproduce and explore phase diagrams obtained through DFT calculations; this indicates the validity of PFP across a wide range of crystal structures and element combinations. This study, which integrates a universal NNP with a GA-based CSP method, highlights the promise of these methods in materials discovery.
preprint: APS/123-QED
I Introduction
Computational atomistic simulations based on a quantum-mechanical description enhances our understanding of materials and even contribute to modern materials design Hafner (2008). Crystal structure prediction (CSP), a process to predict stable crystal structures in given systems, is a crucial prerequisite to harness the computational atomic simulations Oganov (2011); Oganov et al. (2019); Wales (2004); Jain et al. (2016). Despite the well-established methodology for analyzing crystal structures from experimental data, CSP is essential for accelerating materials discovery through the ab initio approach.
While CSP plays an integral role in predicting stable crystal structures, it represents a daunting global optimization task owning to the vast energy landscape. Substituting elements from known crystal structure prototypes is a prevailing method for generating candidate structures Hautier et al. (2011). Although the prototype substitution method often gives reasonable results, its coverage falls short, especially in multicomponent systems. Metaheuristic algorithms offer another approach to creating novel crystal structures. These include methods such as random structure search Pickard and Needs (2011), basin hopping Wales and Doye (1997), minima hopping Amsler and Goedecker (2010); Krummenacher et al. (2024), genetic algorithm (GA) Deaven and Ho (1995); Oganov and Glass (2006); Glass et al. (2006); Lonie and Zurek (2011); Falls et al. (2021), particle swarm optimization Wang et al. (2010), and Bayesian optimization Yamashita et al. (2018); Bisbo and Hammer (2020). Typically, these CSP methods are combined with density functional theory (DFT) calculations to evaluate the formation energy of the candidate structures. However, the time-consuming DFT calculations, which limit the exploration of structures, significantly hinder the efficiency of CSP.
Recent advancements in machine learning potential, interatomic potential fitted by DFT calculations, enable an efficient approach to CSP thanks to their fast and accurate energy evaluation Behler and Parrinello (2007); Bartók et al. (2010); Shapeev (2016); Drautz (2019); Podryabinkin et al. (2019). Because we are often interested in multicomponent systems like ternary or quaternary systems in CSP, it is preferable for the machine learning potential to exhibit scalability with the number of elements and transferability across various systems. Also, high accuracy is required to capture subtle energy differences among distinct structures in CSP. A universal neural network potential (NNP) trained with extensive datasets has been gaining significant attention owing to its scalability and transferability, which includes M3GNet Chen et al. (2019); Chen and Ong (2022), ALIGNN-FF Choudhary et al. (2023), CHGNet Deng et al. (2023), MACE-MP-0 Batatia et al. (2022, 2023), GNoME Merchant et al. (2023), Orb Neumann et al. (2024), EquiformerV2-OMat24 Barroso-Luque et al. (2024); Schmidt et al. (2023), MatterSim Yang et al. (2024), and PFP Takamoto et al. (2022a, b); Jacobs et al. (2025).
Applying a universal NNP to CSP is straightforward. For example, GN-OA Cheng et al. (2022) combines the existing metaheuristic algorithms with a universal NNP. GNoME Merchant et al. (2023) has unveiled various previously unreported stable structures with an extension of prototype substitution and random structure generation. Nevertheless, further development of CSP algorithms employing universal NNPs have room for enhancement. While the existing ab initio CSP methods often implicitly assume the number of energy evaluations to be limited due to the computational cost of DFT, the universal NNPs can allow us the vast number of energy evaluations and consider an entire convex hull concurrently, thereby appreciating the diversity of crystal structures all at once.
Here, we propose a novel CSP method using a universal NNP and GA. We have employed our developed universal NNP called PFP Takamoto et al. (2022b). The current GA is inspired by the sophisticated multi-objective optimization algorithms Deb et al. (2002); Deb and Jain (2014), which enables us to explore the entire composition space by exploiting the low-energy structure regions from optimization history. In this study, we use symmetry-aware random structure generation for the initial population and one of the mutation operations. Nevertheless, we mention that the present GA can be combined with other structure generation methods, such as generative models Xie et al. (2022); Jiao et al. (2023); Zeni et al. (2025), which will be a future work.
The remainder of this paper is organized as follows. Section II describes PFP and a procedure for calculating formation energy with PFP in this study. Section III provides the details of the present GA for CSP. Section IV shows the application of the present GA to a chemically diverse range of elements, from binary to octonary systems.
II Formation energy calculation with PFP
II.1 PFP
In this study, we used our developed universal NNP called the PFP version 6.0.0 Takamoto et al. (2022b), which was trained on around 42 million structures. The PFP can be applied to arbitrary combinations of 72 supported elements (all elements from H to Bi except Tc, Pm, Eu, Tb, Dy, Ho, Er, Tm, Yb, Lu, and Tl). Its training dataset was collected with DFT calculations using plane-wave basis sets and the projector augmented wave (PAW) Blöchl (1994); Kresse and Joubert (1999) with the Perdew–Burke–Ernzerhof (PBE) exchange-correlation functional Perdew et al. (1996) implemented in the Vienna ab-initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996, 1996). Details of the DFT calculations are described in Ref. Takamoto et al., 2022b.
For systems containing V, Cr, Mn, Fe, Co, Ni, Cu, Mo, or W, we prepared the DFT calculations with the generalized gradient approximation (GGA) with PBE functional and GGA with Hubbard $U$ corrections (GGA+U) introduced by Dudarev et al. Dudarev et al. (1998). The $U-J$ parameters in GGA+U were adopted from the Materials Project (MP) Jain et al. (2013) except for Cu and from Wang et al. Wang et al. (2006) for Cu. We trained PFP simultaneously for the DFT calculations with and without Hubbard $U$ corrections. The PFP has two modes to predict energies with and without Hubbard $U$ corrections. We used PFP trained with Hubbard $U$ corrections for oxides and fluorides containing V, Cr, Mn, Fe, Co, Ni, Cu, Mo, or W, and otherwise without Hubbard $U$ corrections. This choice is consistent with MP except for Cu.
II.2 Reference simple systems
For formation energy calculations, we selected reference crystal structures for the 72 supported elements and 2 partially supported elements (Tc and Tl), as listed in Table 1. For each element, we collected crystal structures from MP and searched for the lowest-energy structure after structural relaxation with PFP. The structural relaxation of atomic positions and lattice constants was performed until residual forces were less than $5\times 10^{-3}$ eV/Å. We matched the reference structures with the AFLOW standard encyclopedia of crystallographic prototypes Mehl et al. (2017) using pymatgen Ong et al. (2013). The third column of Table 1 displays the matched prototype structures within 0.1 meV/atom from the lowest-energy structure, which captures the diversity of stable prototype structures across various elements. The fourth column shows the PFP energy differences in meV/atom between two structures: one is the energy-lowest structure in PFP and the other is the energy-lowest structure in DFT calculation in MP. While the energy differences are particularly noticeable for layered or molecular crystal structures of B, O, As, and Br, possibly due to convergence issues in structural relaxations, PFP accurately evaluates stable crystal structures up to approximately 10 meV/atom.
Table 1: Reference simple systems for each element in this study. The second column shows initial structures for structural relaxation taken from MP, recorded with material_id. The third column shows prototype structures within 0.1 meV/atom from the lowest-energy structure. When a prototype is not found, material_id is denoted instead. The fourth column shows the PFP energy differences (meV/atom) between the present reference structures and the on-the-hull structures in MP.
H mp-973783 $\beta$-O, mp-731827 4.3 Sr mp-867202 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0
He mp-614456 A1 (fcc), A3 (hcp) 0.8 Y mp-112 A3 (hcp) 0.0
Li mp-135 A2 (bcc) 0.7 Zr mp-8635 A1 (fcc) 2.2
Be mp-87 A3 (hcp) 0.0 Nb mp-75 A2 (bcc) 0.0
B mp-161 $\beta$-B 40.5 Mo mp-129 A2 (bcc) 0.0
C mp-568286 mp-568286, mp-990424 0.6 Tc mp-113 A3 (hcp) 0.0
N mp-570747 $\gamma$-N 5.7 Ru mp-33 A3 (hcp) 0.0
O mp-1180036 mp-1180036 41.3 Rh mp-74 A1 (fcc) 0.0
F mp-561203 A14 (molecular iodine) 0.0 Pd mp-2 A1 (fcc) 0.0
Ne mp-111 A1 (fcc) 0.0 Ag mp-10597 A3 (hcp) 0.6
Na mp-974920 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0 Cd mp-94 A3 (hcp) 0.0
Mg mp-1056702 A1 (fcc) 1.1 In mp-973111 A3 (hcp) 1.9
Al mp-134 A1 (fcc) 0.0 Sn mp-117 A4 (diamond) 0.0
Si mp-165 Lonsdaleite 5.9 Sb mp-104 A7 ($\alpha$-As) 0.0
P mp-1198724 mp-1198724 9.6 Te mp-567313 A8 ($\gamma$-Se) 0.0
S mp-666931 mp-666931 0.7 I mp-23153 A14 (molecular iodine) 0.0
Cl mp-1008394 A14 (molecular iodine) 0.0 Xe mp-570510 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0
Ar mp-568145 A1 (fcc), A3 (hcp) 0.0 Cs mp-639727 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La) 0.0
K mp-1184764 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0 Ba mp-1096840 A2 (bcc) 0.0
Ca mp-1183455 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La) 0.0 La mp-156 A1 (fcc), A3’ ($\alpha$-La) 0.0
Sc mp-67 A3 (hcp) 0.0 Ce mp-64 A20 ($\alpha$-U) 5.4
Ti mp-46 A3 (hcp) 16.5 Pr mp-38 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0
V mp-146 A2 (bcc) 0.0 Nd mp-123 A1 (fcc), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0
Cr mp-90 A2 (bcc) 0.0 Sm mp-21377 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0
Mn mp-35 A12 ($\alpha$-Mn) 0.0 Gd mp-155 A3 (hcp), C19 ($\alpha$-Sm) 0.0
Fe mp-1271068 A2 (bcc) 0.0 Hf mp-103 A3 (hcp) 0.0
Co mp-54 A3 (hcp) 6.0 Ta mp-50 A2 (bcc) 29.6
Ni mp-23 A1 (fcc) 0.0 W mp-91 A2 (bcc) 0.0
Cu mp-30 A1 (fcc) 0.0 Re mp-8 A3 (hcp) 23.6
Zn mp-1187812 A3 (hcp), C19 ($\alpha$-Sm) 0.1 Os mp-49 A3 (hcp) 0.0
Ga mp-142 A11 ($\alpha$-Ga) 0.0 Ir mp-101 A1 (fcc) 0.0
Ge mp-1007760 Lonsdaleite 4.1 Pt mp-126 A1 (fcc) 0.0
As mp-11 A7 ($\alpha$-As) 23.0 Au mp-1008634 A3 (hcp) 2.0
Se mp-542605 mp-542605 1.1 Hg mp-1184554 A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 2.4
Br mp-1120813 mp-1120813 10.3 Tl mp-151 A1 (fcc) 0.3
Kr mp-976347 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 0.0 Pb mp-20745 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La) 0.1
Rb mp-12628 A1 (fcc), A3 (hcp), A3’ ($\alpha$-La), C19 ($\alpha$-Sm) 12.8 Bi mp-23152 A7 ($\alpha$-As) 0.0
II.3 MP compatibility and energy corrections
Because we used almost the same DFT settings as MP, we applied anion and GGA/GGA+U mixing scheme corrections Wang et al. (2021) into PFP total energies for better predictions across diverse chemical systems. The policy regarding which modes of PFP to use with and without Hubbard $U$ corrections in Section II.1 is compatible with the GGA/GGA+U mixing scheme in MP. The exception is an oxide with Cu, in which case we use a correction for Cu in Ref. Jain et al., 2011. Consequently, we can directly compare a convex hull with PFP and MP for all systems except oxides with Cu.
III GA-based CSP method
In this study, we adopted a common terminology in the field of GA; each cycle for generating new structures from the previous ones is called a generation. A set of structures in a generation is called a population. Likewise, we refer to the process of generating a new structure and evaluating its formation energy as a trial. The present algorithm takes the following inputs: (i) elements (e.g., O–Sr–Ti), (ii) maximum number of atoms in the unit cell, (iii) population size, and (iv) number of total trials. Then, it outputs crystal structures on formation energy convex hulls based on PFP.
The workflow to search for stable crystal structures using GA is illustrated in Fig. 1. First, an initial population of structures is generated by random generation (Section III.1). In each generation, structures are retrieved from the current population, and the promising ones are filtered (Section III.2). Then, an elite population is generated and selected from the promising ones for crossover and mutation (Section III.3). For the elite population, the next population is generated by crossover and mutation (Section III.4). For the next population, the formation energy is calculated after structural relaxation with PFP, as described in Section II.
Note that we present the workflow in Fig. 1 with clarity, assuming a synchronous search is performed by each generation. However, an actual GA implementation can leverage asynchronous search over each trial, which enables a high parallelism. When a structure is retrieved in a trial, its parent population is looked up from its generation. This asynchronous search can be performed effectively on top of the modern optimization framework Optuna Akiba et al. (2019).
Generate initial population (Section III.1)Retrieve structures from current populationFilter promising structures (Section III.2)Generate and select elite population (Section III.3)Generate next population by crossover and mutation (Section III.4)Calculate formation energy after structural relaxationEnough trials?EndYesNo
Figure 1: Schematic diagram of the present GA-based CSP method.
III.1 Random structure generation
To generate the initial population, we follow a random structure generation in Ref. Lyakhov et al., 2010. We split the cell into multiple sub-cells, randomly place atoms in only one of these sub-cells, and then replicated the sub-cell throughout the entire cell. This technique generates a diverse set of structures, as opposed to randomly placing atoms in the entire cell which often generates randomly packed structures. We use symmetry-aware random structure generation by PyXtal Fredericks et al. (2021) for the random structure generation in the sub-cells, thereby improving the search efficiency because stable crystal structures tend to have high symmetry Pickard and Needs (2011).
The random structure generation by PyXtal requires a composition and a space-group type as input. We sampled the number of atoms in the unit cell uniformly up to the given maximum number of atoms in the unit cell and selected a composition from a multinomial distribution where each element has an equal probability. We sampled a space-group type uniformly from all space-group types other than $P1$. As some combinations of a composition and a space-group type are infeasible to generate a crystal structure, we retried the entire sampling process until a crystal structure is successfully generated.
III.2 Population filtering
In this step, we selected structures close to the convex hull with the most recently updated composition. Our algorithm design is based on two intuitions: (1) lower-energy structures should exist near the structures close to the convex hull, and (2) there should be some room for improvement in the neighbors of the compositions that observed recent updates on their lowest energy. Population filtering is similar to the aging evolution strategy Real et al. (2019) in that it ignores the structures that have not been updated for a long time.
Figure 2: The relationship between the parent population, elite population, filtered population, next elite population, and the next population.
Let $M$ be the number of elements in the search, $P_{n}$ be the parent generation, $Q_{n}$ be the elite population used to select the parent generation, and $R_{n}=P_{n}\cup Q_{n}$, as shown in Fig. 2. Let $E(i)$ be the energy of the structure $i$, $r(i)\in[0,1]^{M}$ be the composition of the structure $i$, and $j_{*}(i)$ be the structure with the smallest energy in the neighborhood of the structure $i$ defined as follows:
$j_{*}(i)=\mathop{\mathrm{argmin}}_{j\in\bigcup_{k=1}^{n}P_{k}}\{E(j)\mid\|r(i),r(j)\|_{1}<\delta\},$ (1)
where $\|x\|_{1}=\sum_{k=1}^{d}|x_{i}|$ is the $1$-norm, and $\delta>0$ is the threshold of the $1$-norm of the composition. Here, we use all the structures we have explored so far $\bigcup_{k=1}^{n}P_{k}$ as neighbors instead of $R_{n}$ to maintain the diversity of the elite population. We define the distance to the convex hull of the structure $i$ as
$E(i)-E(j_{*}(i)).$ (2)
Let $n^{*}(i)$ be the generation in which the structure $j_{*}(i)$ was generated. Let $E_{\max}$ and $E_{\min}$ be the maximum and minimum energies over all the structures explored so far defined as follows:
$E_{\max}=\max\{E(i)\mid i\in\bigcup_{k=1}^{n}P_{k}\},$ (3)
$E_{\min}=\min\{E(i)\mid i\in\bigcup_{k=1}^{n}P_{k}\}.$
We later normalize the distance to the convex hull $E(i)-E(j_{*}(i))$ to be between 00 and $1$ by dividing it by $E_{\max}-E_{\min}$.
The population filtering is a procedure to construct a subset $R^{\prime}_{n}\subset R_{n}$. For each element $i$ in $R_{n}$, we calculated the following quantity by considering the distance to the convex hull $E(i)-E(j_{*}(i))$ and the generation difference $n-n^{*}(i)$,
$D(i)=\frac{E(i)-E(j_{*}(i))}{E_{\max}-E_{\min}}\times\alpha^{n-n^{*}(i)},$ (4)
where $\alpha>1$ is a tuning parameter. The parameter $\alpha$ controls the importance of the generation difference. The smaller $D(i)$ indicates that the structure is closer to the convex hull, and composition has been updated more recently. We adopt only the structures $i$ that satisfy $D(i)<\epsilon$, where $\epsilon$ is a tuning parameter. The parameter $\epsilon$ controls the threshold of the distance to the convex hull and the generation difference. Finally, we define the following set $R^{\prime}_{n}$ as the output of this step,
$R^{\prime}_{n}=\left\{i\in R_{n}\mid D(i)<\epsilon\right\}.$ (5)
III.3 Elite population selection
The input of the elite population selection is the set $R^{\prime}_{n}$ of structures selected by the population filtering, and the output is the elite population $Q_{n+1}$ to generate the next generation. The elite population is a set of structures to generate the next generation, and we select structures to apply crossover and mutation in the subsequent steps. We can reduce computational costs and search for stable structures more efficiently by applying crossover and mutation only to promising structures.
A commonly used method for selecting the elite population involves calculating the fitness of each structure and selecting the structures with the highest fitness Larsen et al. (2017); Vilhelmsen and Hammer (2012, 2014). For instance, ASE Larsen et al. (2017) defines it using the reciprocal of the energy of each structure. It only selects the structure with the highest fitness among the structures with the same composition. However, it does not use quantitative information about the composition and cannot maintain the diversity of the composition in the elite population.
Conversely, we selected the elite population using quantitative information about the composition in addition to the energy of the structure since the convex hull depends on both the energy and the composition. We found that the convex hull in the energy–composition space can be identified with the Pareto front in the multi-objective optimization problem (Fig. 3). The convex hull is designed to spread as much as possible downward in the energy axis direction, and the Pareto front is designed to spread as much as possible downward in the objective variable space.
Figure 3: The similarity between the convex hull in the energy–composition space and the Pareto front in the multi-objective optimization problem.
We apply the elite population selection method used in the well-known genetic algorithms for multi-objective optimization problems, such as NSGA-II Deb et al. (2002) and NSGA-III Deb and Jain (2014). Notably, NSGA-II and NSGA-III are powerful algorithms for multi-objective optimization and are implemented in various black-box optimization software such as Refs. Akiba et al. (2019) and Blank and Deb (2020). The difference between NSGA-II and NSGA-III is the procedure for tie-breaking. In NSGA-II, a procedure is used to select structures with the smallest “crowding distance” for a set of structures of a certain rank. In NSGA-III, a procedure is used to draw a line called a “reference line” for a set of structures of a certain rank, calculate the crowding degree for each line, and select structures in order of less crowded. In general, NSGA-II is effective when the number of objective variables is 2–3, and NSGA-III is effective when it is 4 or more. The number of elements in the system is equal to the dimension of the multi-objective optimization in our problem.
Our proposed elite population selection method is summarized as follows. Note that we use the same notations as in Section III.2.
  1. 1.
    Initialize the elite population $Q_{n+1}$ as an empty set.
  2. 2.
    Sort the set $R^{\prime}_{n}$ by non-dominated sort with a partial order relation $\prec$ that depends on both energy and composition, and assign a rank to each structure $i\in R^{\prime}_{n}$.
  3. 3.
    Until the elite population $Q_{n+1}$ exceeds the specified size, select structures in ascending order of rank and add them to the elite population $Q_{n+1}$.
  4. 4.
    When all structures of a certain rank $k$ are added to the elite population $Q_{n+1}$, if the size of the elite population $Q_{n+1}$ exceeds the specified size, select structures to add to the elite population $Q_{n+1}$ from the structures of rank $k$ according to a specific procedure.
Our proposed method is described in the following sections by defining the partial order $\prec$ and the tie-breaking procedure for ranks. For later applications in Section IV, we use an NSGA-III-like tie-breaking procedure described in Section III.3.3. We observed that the NSGA-III-like tie-breaking procedure outperformed the NSGA-II-like one in our preliminary experiments, consistent with the general trend that NSGA-III is more effective in higher-dimensional optimization problems.
III.3.1 Partial order used for non-dominated sort
We formally define the energy–composition space as a subset of $\mathbb{R}^{M+1}$,
=$\mathcal{R}=\left\{(r_{1},\ldots,r_{M},E)\in[0,1]^{M}\times\mathbb{R}\mathrel{}\middle|\mathrel{}\sum_{a=1}^{M}r_{a}=1\right\}.$ (6)
For each $a\in\{1,\ldots,M\}$, let $r(i)_{a}$ be the ratio of the $a$-th element in the composition $r(i)$. We define the partial order $\prec$ using the convex hull in the energy–composition space. Let $H^{0}$ be the convex hull of the structures in $R^{\prime}_{n}$ in $\mathcal{R}$. Let $H^{1}$ be the convex hull of the structures in $R^{\prime}_{n}\setminus H^{0}$, and $H^{2}$ be the convex hull of the structures in $R^{\prime}_{n}\setminus(H^{0}\cup H^{1})$. Repeat this process, and let $H^{m}$ be the convex hull of the structures in $R^{\prime}_{n}\setminus\left(\bigcup_{k=0}^{m-1}H^{k}\right)$. By peeling off the thin skin from $R^{\prime}_{n}$ to obtain sets $H^{0},H^{1},H^{2},\ldots$, we define the partial order $\prec$ like the relationship of $i\prec j$ holds if and only if $i\in H^{k}$ and $j\in H^{l}$ where $k<l$. We can assign a rank to each structure by sorting $R^{\prime}_{n}$ using the partial order $\prec$. We denote the rank assigned to each structure $i$ as $k_{i}$.
III.3.2 Tie-breaking procedure in NSGA-II-like CSP algorithm
First, we define a procedure using the crowding distance similar to NSGA-II. The definition of crowding distance is different because we consider a different problem from the usual multi-objective optimization problem.
Let $H^{(k,a)}$ be the set of $H^{k}$, the output of the population filtering with rank $k$, sorted by $r(i)_{a}$. For each $a\in\{1,\ldots,M\}$ and structure $i$, let $r(i)_{a,\mathrm{left}}$ and $r(i)_{a,\mathrm{right}}$ be the ratio of the $a$-th element in the structures on both sides of $i$ in $H^{(k,a)}$, and let $\Delta_{r}(i,a)$ be the difference between them,
$\Delta_{r}(i,a)=r(i)_{a,\mathrm{right}}-r(i)_{a,\mathrm{left}}.$ (7)
Similarly, for the energy $E(i)$, let $H^{(k,E)}$ be the set of $H^{k}$ sorted by $E(i)$. For each structure $i$, let $E(i)_{\mathrm{right}}$ and $E(i)_{\mathrm{left}}$ be energies of the structures on both sides of $i$ in $H^{(k,E)}$ and $\Delta_{E}(i)$ be the difference of them as follows:
$\Delta_{E}(i)=E(i)_{\mathrm{right}}-E(i)_{\mathrm{left}}.$ (8)
Note that the energy of each structure is min-max normalized for stability.
Using the $\Delta_{r}(i,a)$ and $\Delta_{E}(i)$ defined above, we define the crowding distance $c(i)$ for the structures $i$ in $H^{k}$ as follows:
$c(i)=\Delta_{E}(i)\times\sum_{a=1}^{M}\Delta_{r}(i,a).$ (9)
We can order structures with the same rank by selecting structures with small crowding distance $c(i)$. The algorithm is summarized in Algorithm 1.
Algorithm 1 Selection of $\ell$ Structures with Smallest Crowding Distance (NSGA-II Inspired)
1:Population set $H^{k}$ with rank $k$, number of selected structures $\ell$
2:Set $\hat{H}^{k}$ containing $\ell$ structures with smallest $c(i)$
3:for each $a\in\{1,\ldots,M\}$ do
4:     Sort $H^{k}$ by $r(i)_{a}$ to create $H^{(k,a)}$
5:     for each structure $i\in H^{k}$ do
6:         Identify neighboring structures in $H^{(k,a)}$
7:         Compute $r(i)_{a,\mathrm{left}}$ and $r(i)_{a,\mathrm{right}}$
8:         Compute $\Delta_{r}(i,a)=r(i)_{a,\mathrm{right}}-r(i)_{a,\mathrm{left}}$
9:     end for
10:end for
11:Sort $H^{k}$ by energy $E(i)$ to create $H^{(k,E)}$
12:for each structure $i\in H^{k}$ do
13:     Identify neighboring structures in $H^{(k,E)}$
14:     Compute $E(i)_{\mathrm{left}}$ and $E(i)_{\mathrm{right}}$
15:     Compute $\Delta_{E}(i)=E(i)_{\mathrm{right}}-E(i)_{\mathrm{left}}$
16:end for
17:for each structure $i\in H^{k}$ do
18:     Compute crowding distance:
$c(i)=\Delta_{E}(i)\times\sum_{a=1}^{M}\Delta_{r}(i,a)$
19:end for
20:Sort structures in $H^{k}$ by $c(i)$ in ascending order
21:Select the first $\ell$ structures with the smallest $c(i)$ to form $\hat{H}^{k}$
22:return $\hat{H}^{k}$
III.3.3 Tie-breaking procedure in NSGA-III-like CSP algorithm
Next, we define a procedure using a reference line similar to NSGA-III. The procedure for tie-breaking using the reference line in NSGA-III is performed as follows:
  1. 1.
    Normalize the value of each objective variable to be non-negative.
  2. 2.
    In the space of the objective variables, draw a line connecting the origin and the predetermined multiple reference points and construct multiple reference lines.
  3. 3.
    For each structure, project its objective variable values onto each reference line, select the reference line with the smallest distance to the projection point, and add it to the neighborhood set of each reference line.
  4. 4.
    Select the reference line with the smallest neighborhood set, select the structure closest to the reference line in the neighborhood set, and add this to the elite population by excluding it from the neighborhood set.
  5. 5.
    Repeat the above procedure until the elite population reaches the specified size.
To clarify the procedure, we need to define reference points and how to measure distances. We adopt the following distance:
$d(i,j)=|E(i)-E(j)|\times\sum_{a=1}^{M}|r(i)_{a}-r(j)_{a}|.$ (10)
We constructed the reference points as follows. For each composition axis $a\in\{1,\ldots,M\}$, select the structure with the maximum $r(i)_{a}$ in $H^{k}$ and call them $i_{1},\ldots,i_{M}$. Then, consider an $(M-1)$-dimensional simplex with $i_{1},\ldots,i_{M}$ as vertices, and place points $r_{1},\ldots,r_{K}$ uniformly at regular intervals on it, and use these as reference points. Here, $K$ is determined by the number of points to be placed on the edge of the $(M-1)$-dimensional simplex, and if $k+1$ points are placed on the edge, $K=\binom{M+k-1}{k}$. The algorithm is summarized in Algorithm 2.
1:Population set $H^{k}$ with rank $k$, number of selected structures $\ell$
2:Set $\hat{H}^{k}$ containing $\ell$ structures distributed along the reference lines
3:Normalize each objective variable to be non-negative
4:Construct reference lines:
5:Define $K$ reference points by selecting structures $i_{1},\ldots,i_{M}$ with maximum $r(i)_{a}$ for each axis $a\in\{1,\ldots,M\}$
6:Place $K$ points $r_{1},\ldots,r_{K}$ uniformly on an $(M-1)$-dimensional simplex
7:for each structure $i\in H^{k}$ do
8:     for each reference line $j$ do
9:         Project $i$ onto reference line $j$
10:         Compute distance $d(i,j)=|E(i)-E(j)|\times\sum_{a=1}^{M}|r(i)_{a}-r(j)_{a}|$
11:     end for
12:     Assign $i$ to the reference line $j^{*}$ with the smallest $d(i,j)$
13:     Add $i$ to the neighborhood set of $j^{*}$
14:end for
15:while size of $\hat{H}^{k}$ is less than $\ell$ do
16:     Select reference line $j^{*}$ with the smallest neighborhood set
17:     Select structure $i^{*}$ in the neighborhood set of $j^{*}$ closest to the reference line
18:     Add $i^{*}$ to $\hat{H}^{k}$ and remove it from the neighborhood set of $j^{*}$
19:end while
20:return $\hat{H}^{k}$
Algorithm 2 Selection of $\ell$ Structures based on the Reference Line (NSGA-III Inspired)
III.4 Crossover and mutation
Crossover and mutation are the processes of generating new structures $P_{n+1}$ from the elite population $Q_{n+1}$. Crossover involves combining two parent structures to generate a child structure, and mutation involves changing the structure of a parent structure.
The crossover and mutations used in our proposed method are listed in Table 2. Several methods are implemented in ASE Larsen et al. (2017), and others are our original methods. To customize the methods in ASE, we modify the implementation and release them as a package.  [^1] The modified crossover and mutation methods are available at https://pypi.org/project/pfn-ase-extras/
Table 2: Crossover and mutation methods used in the proposed GA-based CSP method.
Method Description Implementation
Selection
probability
Cut-and-splice crossover
Generate a child structure by cutting two parent structures
at their unit cells along a crystal plane and joining them Vilhelmsen and Hammer (2012).
Modified ASE $1/2$
Random composition mutation Generate a child structure by randomly changing its composition. ASE $1/12$
Strain mutation
Distort the unit cell by applying a random strain matrix
to the basis vectors.
ASE $1/12$
Rattle mutation Perturb the atomic positions by a random displacement. ASE $1/12$
Permutation mutation Swap the atomic positions of two selected atoms. ASE $1/12$
Random atom deletion mutation
Generate a child structure by randomly removing one atom from
the parent structure.
Original $1/12$
Random structure generation mutation Generate a structure by the random structure generation in Section III.1. Original $1/12$
The cut-and-splice crossover is a method to generate a child structure by cutting two parent structures at their unit cells along a crystal plane and joining them. Originally, its implementation in ASE only allowed crossover between parent structures with the same composition. We modify the implementation to allow crossover between parent structures with different compositions.
The random atom deletion mutation is a method to generate a child structure by randomly removing one atom from the parent structure. This mutation is effective to change the composition of the structure and explore the composition space.
The random structure generation mutation is a method to generate a child structure by the random structure generation in Section III.1. This mutation can effectively maintain the diversity of the composition in the population.
IV Results and discussion
In this section, we present the results of our proposed method for CSP using PFP and GA. First, we evaluated the performance of the proposed method by comparing it with a random search in Section IV.1. We then demonstrated the effectiveness of the proposed method by performing CSP searches for binary and ternary systems with MP in Section IV.2. The results demonstrate that the proposed method can efficiently explore the search space and discover new crystal structure candidates that update the convex hull of MP. For the energy evaluation, we used the energy corrections described in Section II.3.
IV.1 Quantitative comparison with random search and MP
Figure 4: Comparison between random search (orange) and the proposed method (blue) by the normalized hull volume history. The hatchings indicate the minimum and maximum normalized hull volumes within the five searches. The black dashed line indicates the normalized volumes for the convex hull constructed from MP structures.
We quantitatively evaluated the proposed method compared with a random search. The random search was performed using the method described in Section III.1. In this experiment, we fixed the following inputs and parameters: the maximum number of atoms in the unit cell was 128, the population size was 250, the total number of trials was 50,000, and $\alpha=2.4$, $\epsilon=0.25$. The genetic operations and their selection probabilities are listed in Table 2. We chose four systems ranging from ternary to octonary systems: O–Sr–Ti and Ba–Ca–Cu–O–Tl as well known oxide systems Zeni et al. (2025), and Cu–Ga–Rh–Sc and Co–Cr–Cu–Fe–Mn–Ni–Ti–V as random alloy systems. The Hubbard $U$ correction for Cu is applied for our method, while it is not applied for MP, which may affect the comparison in the Ba–Ca–Cu–O–Tl system, as mentioned in Section II.1.
We evaluated each search by the volume of the convex hull, which can assess both the depth of the convex hull and the diversity of the included structures Donaldson et al. (2024). The hull volume value of the $M$-element system is normalized by the unit hull volume $V_{{\rm unit},M}=1\,\mathrm{eV}/M!$, which corresponds to the volume of the convex hull with only one crystal structure on the hull with formation energy of -$\mathord{-}1$ eV/atom. We performed the search by five times for each system because the CSP search involves randomness. Additionally, we compared the normalized hull volume with that of MP. We performed structure optimization by PFP for MP structures within 0.2 eV/atom in MP convex hull to construct the MP convex hull. The absolute value of the normalized hull volume depends on the element system and cannot be compared across different systems.
Figure 4 illustrates how the normalized hull volume changed as the search progressed for each system, indicating that the proposed method outperformed random search for all the element systems. The final mean normalized hull volumes for each system are listed in Table 3. Here, we denote the mean normalized hull volume as $V_{\mathrm{mean}}$, the maximum as $V_{\mathrm{max}}$, and the minimum as $V_{\mathrm{min}}$. The improvements of final mean hull volumes of the proposed method compared with random search are 0.005, 0.011, 0.022, and 0.022 for O–Sr–Ti, Cu–Ga–Rh–Sc, Ba–Ca–Cu–O–Tl, and Co–Cr–Cu–Fe–Mn–Ni–Ti–V respectively. These improvements correspond to updates of the convex hull by 5 to 22 meV/atom if only one entry exists on the convex hull. We mention that the conversion of a normalized hull volume to a formation energy update depends on the shape of the convex hull. The updated hull volume by the proposed method compared with that of MP is 0.015, 0.038, 0.070, and -0.002 for O–Sr–Ti, Cu–Ga–Rh–Sc, Ba–Ca–Cu–O–Tl, and Co–Cr–Cu–Fe–Mn–Ni–Ti–V, respectively; this corresponds to 0.3, 1.9, 1.3, and -0.1% of MP hull volume, respectively. The proposed method achieved larger convex hull volumes than MP for systems except for the Co–Cr–Cu–Fe–Mn–Ni–Ti–V. In the five searches of the octonary alloy system, there are two crystal structures that are never found by proposed methods, as discussed later; this is attributed to the smaller convex hull volume compared to MP.
Figure 4 also depicts the efficiency of the CSP search with the proposed method. The number of generations required to reach 99% of each final hull volume for O–Sr–Ti, Cu–Ga–Rh–Sc, Ba–Ca–Cu–O–Tl, and Co–Cr–Cu–Fe–Mn–Ni–Ti–V were 7, 47, 16, and 61, respectively, for the proposed method, and 4, 66, 36, and 115, respectively, for a random search; this indicates that the proposed method can efficiently explore the search space, especially in multicomponent systems. Additionally, the figure exhibits that the proposed method has smaller fluctuations in the hull volume compared to random search. The fluctuation of the final hull volume for the five searches is calculated as $(V_{\rm{max}}-V_{\rm{min}})/V_{\rm{mean}}$, which is 0.05, 0.35, 0.08, and 0.72% for the proposed method and 0.20, 0.44, 1.14, and 3.29% for random search for O–Sr–Ti, Cu–Ga–Rh–Sc, Ba–Ca–Cu–O–Tl, and Co–Cr–Cu–Fe–Mn–Ni–Ti–V, respectively; this indicates that the proposed method can stably explore the search space.
Table 3: Mean normalized hull volume and average number of MP hull structures compared with our CSP hull. The second columns show the mean normalized hull volume of the proposed method, while the last two columns show that for the random search and the normalized hull volume of MP, respectively. The third to fifth columns show the average number of MP structures below, around, and above the CSP hull by the proposed method, respectively (See the main text for the detailed definition). The sixth column shows the average ratio of the number of MP structures around or above the CSP hull.
System $V_{\mathrm{proposed}}$ Below hull Around hull Above hull Reproducibility $V_{\mathrm{random}}$ $V_{\mathrm{MP}}$
O–Sr–Ti 5.056 0 10.2 3.8 1.00 5.051 5.041
Cu–Ga–Rh–Sc 2.056 4.8 13.2 5 0.79 2.045 2.018
Ba–Ca–Cu–O–Tl 5.687 10.4 15.8 2.8 0.64 5.666 5.617
Co–Cr–Cu–Fe–
Mn–Ni–Ti–V
1.601 8.6 38.2 7.2 0.84 1.579 1.603
Figure 5: Distribution of the energy above the hull of the crystal structures of MP. The convex hull is reconstructed using the CSP results obtained using the method proposed in Section IV.1. The results of the five searches are aggregated and presented by the system.
We validated the effectiveness of our proposed method for actual applications by comparing the convex hulls of CSP and that of MP. Table 3 shows the average number of MP structures below, around, and above the convex hull of CSP searches. The structures around the hull are defined as those with energy above hull within 0.01 eV/atom. The structures below and above the hull are those with energy above hull less than -0.01 eV/atom and greater than 0.01 eV/atom, respectively. The average ratio of the number of MP structures around and above the convex hull to the total number of structures is 1.00, 0.79, 0.63, and 0.84 for O–Sr–Ti, Cu–Ga–Rh–Sc, Ba–Ca–Cu–O–Tl, and Co–Cr–Cu–Fe–Mn–Ni–Ti–V, respectively. Although the trends vary by element systems, around 80% of MP structures are reproduced by CSP. In particular, all MP structures are reproduced in the O–Sr–Ti system.
Figure 5 illustrates an energy above hull distribution of MP structures with the CSP hull by the proposed method. The results of the five searches are aggregated for each element system. As demonstrated, many MP structures are concentrated around the CSP hull. In this figure, MP structures with negative energy above hull indicate that the CSP search by the proposed method missed more stable structures in MP. For the octonary system of Co–Cr–Cu–Fe–Mn–Ni–Ti–V, two crystal structures shows energy above hull less than -0.1 eV/atom, Cr2Ti (mp-1425) and Mn2Ti (mp-1949). These structures are not found in all the five CSP searches, while random search was able to find Cr2Ti twice and Mn2Ti once in five searches in the potential energy landscape of PFP.
IV.2 CSP for binary and ternary systems
We demonstrated the capability of our CSP method using PFP to discover new crystal structure candidates for binary and ternary systems. The search conditions were the same as those in Section IV.1; however, the structures from MP are added to the initial population of the search. The crystal structures obtained from the search were re-evaluated using DFT calculations to verify their validity. We report the results for systems In–Li, As–V, Al–Li–Pd, and La–Mo–O shown in Figs. 6 and 7.
Figure 6 depicts the convex hulls by our CSP methods and MP, respectively. DFT calculations with VASP were performed for these structures using the same settings as in Section II.1. By comparing convex hulls, the structures identified through our CSP method are confirmed to be also stable in DFT calculations. These findings suggest that the CSP approach using PFP is highly effective in discovering new crystal structure candidates.
Figure 7 shows comparisons of MP and CSP phase diagrams for ternary systems. Numerous promising new crystal candidates have also been discovered in the ternary systems. Our CSP methodology demonstrates the capability to explore and update the entire phase diagram comprehensively.
Crystal structures that update the convex hull of MP in the systems shown in Figs. 6 and 7 are listed in Table 4 222 These crystal structures in xyz formats are available at Supplemental Materials. . The space-group type was identified using spglib with symprec = 0.01 and angle_tolerance = 5 Togo et al. (2024). Additionally, we matched these crystal structures with AFLOW prototypes Mehl et al. (2017); Hicks et al. (2019, 2021); Eckert et al. (2024) by assigning the AFLOW label using aviary.wren Goodall et al. (2022). Several crystal structures were not found as prototype structures in the AFLOW database, indicating that our CSP method can discover novel crystal structures. Some of the listed crystal structures are visualized in Fig. 8.
Figure 6: Binary phase diagrams of (a) In–Li and (b) As–V systems. Red and blue lines show the convex hulls of the formation energy of our CSP results and MP, respectively. The energy is evaluated by VASP for solid lines, while PFP is used for dashed lines.
Figure 7: Phase diagrams of ternary systems. Black and blue lines show the convex hulls of MP and our CSP results, respectively. Here, formation energies were evaluated with DFT calculations. Blue and red hatchings indicate the differences between the MP and our CSP convex hulls. The blue region shows the CSP results updating the MP convex hull. Markers display the simplices of the CSP convex hull, while orange squares indicate those with energy updates greater than 10 meV/atom.
Table 4: List of crystal structures that update the convex hull. The fourth column shows the matched AFLOW prototype if it exists.
Formula Space-group type
Hull energy update
$[\mathrm{meV/atom}]$
AFLOW prototype
In2Li2 $Pmma$$\mathord{-}35$ AB_oP4_51_e_f-001
In2Li7 $P\overline{1}$$\mathord{-}1$
In4Li11 $P\overline{1}$$\mathord{-}10$
In4Li9 $C2/m$$\mathord{-}5$
As4V5 $I4/m$$\mathord{-}45$ A4B5_tI18_87_h_ah-001
Al2Li2Pd $C2/m$$\mathord{-}42$
Al2Li2Pd2 $P\overline{3}m1$$\mathord{-}34$
Al2Li2Pd6 $P2_{1}/m$$\mathord{-}33$
Al2Li4Pd2 $P4_{2}/mmc$$\mathord{-}14$
Al2Li4Pd6 $P2_{1}/m$$\mathord{-}25$
Al2Li6Pd8 $Im\overline{3}m$$\mathord{-}19$
Al2Li7 $P\overline{1}$$\mathord{-}4$
Al2LiPd6 $P\overline{1}$$\mathord{-}18$
Al3Pd7 $C2/m$$\mathord{-}4$
Al4Li3Pd2 $P\overline{3}m1$$\mathord{-}22$
Al4LiPd2 $P\overline{1}$$\mathord{-}8$
AlLi5Pd2 $P\overline{4}m2$$\mathord{-}4$
AlLiPd4 $Cmmm$$\mathord{-}33$
AlPd11 $Cmmm$$\mathord{-}3$
AlPd17 $P4/mmm$$\mathord{-}1$
Li33Pd55 $P\overline{1}$$\mathord{-}36$
LiPd3 $Cmmm$$\mathord{-}33$ AB3_oC8_65_a_bf-001
La2O6 $C2/m$$\mathord{-}58$
La2Mo2O6 $R3c$$\mathord{-}1$
La2Mo4O15 $P\overline{1}$$\mathord{-}100$
La4Mo6O24 $P\overline{1}$$\mathord{-}105$
La4MoO9 $C2$$\mathord{-}12$
Mo3O9 $Cm$$\mathord{-}60$
Figure 8: Crystal structures that update the convex hull.
V Conclusion
We have presented the efficient GA-based CSP method using our developed universal NNP, PFP. Our method comprehensively explores the entire composition space and successfully discovers numerous stable crystal structures. The proposed method even identified unregistered stable crystal structures in MP, and subsequent DFT calculations validated their stability. Given that CSP requires precise formation energy accuracy to distinguish various polymorphs, these results also demonstrate the validity of PFP for a wide range of crystal structures and element combinations. To efficiently search the entire composition space, we have proposed the novel elite population selection method inspired from multi-objective optimization techniques, which considers both formation energies and compositions. Since the proposed GA-based CSP method is agnostic to how to generate initial populations, it would be fruitful in future works to combine it with other structure generation methods to tackle the vast search spaces in CSP, such as generative models for crystal structures. Our present method and results show significant promise for accelerating materials discovery.
Acknowledgements.
We would like to thank our colleagues from Preferred Networks, Inc. for helpful discussions and support, including Shuhei Watanabe. PFP v6 was developed using the National Institute of Advanced Industrial Science and Technology’s AI Bridging Cloud Infrastructure (ABCI) in addition to in-house supercomputers in Preferred Networks, Inc.
Competing interests
T.S, H.I, K.N, K.S, C.S, and S.T are employees of Preferred Networks, Inc. Preferred Computational Chemistry Inc., a joint venture between Preferred Networks Inc. and ENEOS Corporation, offers a software-as-a-service called Matlantis, which incorporates a product based on this research.

### v2

@@
 (2024), genetic algorithm (GA) Deaven and Ho (1995); Oganov and Glass (2006); Glass et al.
+(2006); Oganov et al.
+(2011); Lyakhov et al.
 (2013); Lonie and Zurek (2011); Falls et al.
@@
 (2024); Schmidt et al.
-(2023), MatterSim Yang et al.
+(2023a), MatterSim Yang et al.
 (2024), and PFP Takamoto et al.
@@
 (2025).
-Applying a universal NNP to CSP is straightforward.
-For example, GN-OA Cheng et al.
-(2022) combines the existing metaheuristic algorithms with a universal NNP.
-GNoME Merchant et al.
-(2023) has unveiled various previously unreported stable structures with an extension of prototype substitution and random structure generation.
+Applying a machine learning potential to CSP is relatively straightforward and it still achieves a significant improvement in efficiency Jacobsen et al.
+(2018); Podryabinkin et al.
+(2019); Tong et al.
+(2018); Wang et al.
+(2020); Park and Wolverton (2020); Kang et al.
+(2022); Cheng et al.
+(2022); Schmidt et al.
+(2023b); Merchant et al.
+(2023); Chang et al.
+(2024).
 Nevertheless, further development of CSP algorithms employing universal NNPs have room for enhancement.
@@
 (2025), which will be a future work.
+In recent years, research on multi-objective genetic algorithms has been emerging Omee et al.
+(2023, 2025); Donaldson et al.
+(2024).
+Donaldson et al.
+Donaldson et al.
+(2024) proposed a method that combines a multi-objective genetic algorithm with a universal NNP to explore the entire convex hull of a given system.
+Exploring the entire convex hull requires significantly longer simulations than composition-fixed searches.
+However, the existing method quickly fall into local optima and are not able to benefit from prolonged simulations as shown in Section IV.
+We propose a method that allows long-time search without being trapped in local minima by dynamically shifting the search region via the forgetting of outdated structures.
+Moreover, even with relatively long simulations, the search space of crystal structure prediction remains vast.
+Exhaustive search is of course infeasible, and random search is insufficient in terms of sample efficiency.
+To explore efficiently, it is essential to go beyond conventional genetic algorithms and exploit similarities between structures with different compositions.
+We propose a niching strategy that considers both proximity to the convex hull and similarities between formation energy and composition.
+While the idea of emphasizing structures near the convex hull is shared with Donaldson et al.
+Donaldson et al.
+(2024), our formulation was developed independently.
 The remainder of this paper is organized as follows.
@@
 Our algorithm design is based on two intuitions: (1) lower-energy structures should exist near the structures close to the convex hull, and (2) there should be some room for improvement in the neighbors of the compositions that observed recent updates on their lowest energy.
-Population filtering is similar to the aging evolution strategy Real et al.
+Population filtering is similar to the aging evolution strategy Zhu et al.
+(2012); Real et al.
 (2019) in that it ignores the structures that have not been updated for a long time.
@@
 III.4 Crossover and mutation
+Table 2: Crossover and mutation methods used in the proposed GA-based CSP method.
+Method Description
+Selection
+probability
+Cut-and-splice crossover
+Generate a child structure by cutting two parent structures
+at their unit cells along a crystal plane and joining them Oganov et al. (2010, 2011).
+$1/2$
+Random composition mutation Generate a child structure by randomly changing its composition Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
+Strain mutation
+Distort the unit cell by applying a random strain matrix
+to the basis vectors Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013).
+$1/12$
+Rattle mutation Perturb the atomic positions by a random displacement Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
+Permutation mutation Swap the atomic positions of two selected atoms Glass et al. (2006); Oganov et al. (2010); Lyakhov et al. (2013). $1/12$
+Random atom deletion mutation
+Generate a child structure by randomly removing one atom from
+the parent structure.
+$1/12$
+Random structure generation mutation Generate a structure by the random structure generation in Section III.1. $1/12$
 Crossover and mutation are the processes of generating new structures $P_{n+1}$ from the elite population $Q_{n+1}$.
@@
 Several methods are implemented in ASE Larsen et al.
-(2017), and others are our original methods.
-To customize the methods in ASE, we modify the implementation and release them as a package.
-[^1] The modified crossover and mutation methods are available at https://pypi.org/project/pfn-ase-extras/
-Table 2: Crossover and mutation methods used in the proposed GA-based CSP method.
-Method Description Implementation
-Selection
-probability
-Cut-and-splice crossover
-Generate a child structure by cutting two parent structures
-at their unit cells along a crystal plane and joining them Vilhelmsen and Hammer (2012).
-Modified ASE $1/2$
-Random composition mutation Generate a child structure by randomly changing its composition. ASE $1/12$
-Strain mutation
-Distort the unit cell by applying a random strain matrix
-to the basis vectors.
-ASE $1/12$
-Rattle mutation Perturb the atomic positions by a random displacement. ASE $1/12$
-Permutation mutation Swap the atomic positions of two selected atoms. ASE $1/12$
-Random atom deletion mutation
-Generate a child structure by randomly removing one atom from
-the parent structure.
-Original $1/12$
-Random structure generation mutation Generate a structure by the random structure generation in Section III.1. Original $1/12$
-The cut-and-splice crossover is a method to generate a child structure by cutting two parent structures at their unit cells along a crystal plane and joining them.
-Originally, its implementation in ASE only allowed crossover between parent structures with the same composition.
-We modify the implementation to allow crossover between parent structures with different compositions.
+(2017) and we modify some of the implementation 111 The modified crossover and mutation methods are released and available at https://pypi.org/project/pfn-ase-extras/.
+.
+The cut-and-splice crossover is a method to generate a child structure by cutting two parent structures at their unit cells along a crystal plane and joining them Deaven and Ho (1995); Vilhelmsen and Hammer (2012).
+Because its implementation in ASE only allowed crossover between parent structures with the same composition, we modify it to allow crossover between parent structures with different compositions, which is already known in the field of CSP Oganov et al.
+(2010, 2011).
 The random atom deletion mutation is a method to generate a child structure by randomly removing one atom from the parent structure.
@@
 In this section, we present the results of our proposed method for CSP using PFP and GA.
-First, we evaluated the performance of the proposed method by comparing it with a random search in Section IV.1.
-We then demonstrated the effectiveness of the proposed method by performing CSP searches for binary and ternary systems with MP in Section IV.2.
+First, we conducted an ablation study of the population filtering in Section IV.1 to demonstrate the effectiveness of the proposed method.
+Secondly, we evaluated the performance of the proposed method by comparing it with a random search in Section IV.2.
+We then demonstrated the effectiveness of the proposed method by performing CSP searches for binary and ternary systems with MP in Section IV.3.
 The results demonstrate that the proposed method can efficiently explore the search space and discover new crystal structure candidates that update the convex hull of MP.
 For the energy evaluation, we used the energy corrections described in Section II.3.
-IV.1 Quantitative comparison with random search and MP
-Figure 4: Comparison between random search (orange) and the proposed method (blue) by the normalized hull volume history.
+IV.1 Ablation study of population filtering
+Figure 4: Compositional dispersion history of the Br–Pb–Rh system.
+The Average Nearest Neighbor Index (ANNI) is plotted against the generation number.
+The blue line indicates the case with population filtering, while the purple line indicates the case without population filtering.
+The shaded area represents the standard deviation of ANNI values across 10 experiments.
+Lower ANNI values indicate greater compositional skewness.
+Figure 5: Compositional distribution of structures explored at each generation in the Br–Pb–Rh system.
+The upper panel shows the case with population filtering, while the lower panel shows the case without population filtering.
+From left to right, the generations are 5, 25, 50, and 100.
+In this section, we conducted an ablation study of the population filtering.
+Population filtering is performed immediately prior to elite population selection in each generation to preserve the diversity of candidate compositions.
+The ablation study compared two cases: one with population filtering and one without.
+We demonstrated that implementing population filtering enables exploration of a wider range of structures across various compositions compared to its absence.
+In this experiment, we examined the Br–Pb–Rh system with the following parameters: maximum number of atoms in the unit cell was 128, the population size was 250, the total number of trials was 50,000, $\alpha=2.4$, and $\epsilon=0.25$.
+We repeated the search 10 times for each case to account for the stochastic nature of the CSP search.
+The genetic operations and their selection probabilities are specified in Table 2.
+We conducted comparisons under two conditions: with and without enabling population filtering as described in Section III.2.
+To quantitatively assess how compositional dispersion evolves across generations, we employed the Average Nearest Neighbor Index (ANNI) Clark and Evans (1954) to evaluate the compositional homogeneity of structures explored in each generation.
+Lower ANNI values indicate greater compositional skewness.
+As shown in Figure 4, when population filtering is enabled, the ANNI values are significantly higher during early generations compared to when filtering is disabled, indicating greater compositional dispersion.
+However, beginning approximately from the 40th generation onward, ANNI values become comparable regardless of whether population filtering is active or not, suggesting that compositional distribution becomes similarly skewed throughout both search phases.
+Figure 5 illustrates the distribution of compositions of structures explored at each generation for one search.
+At generations 5 and 25, we observe that enabling population filtering results in a more broadly distributed range of structural compositions.
+From generation 50 onward, even when population filtering is active, the distribution of compositions becomes narrower, exhibiting patterns similar to when population filtering is disabled.
+Conversely, when population filtering is disabled, the explored structures show concentration within a narrow compositional range, and this range remains relatively stable across generations.
+This demonstrates that population filtering enables exploration of structures across a wider portion of the search space.
+IV.2 Quantitative comparison with random search and MP
+Figure 6: Comparison between random search (blue) and the proposed method (orange) by the normalized hull volume history.
 The hatchings indicate the minimum and maximum normalized hull volumes within the five searches.
@@
 Figure 7: Distribution of the energy above the hull of the crystal structures of MP.
-The convex hull is reconstructed using the CSP results obtained using the method proposed in Section IV.1.
+The convex hull is reconstructed using the CSP results obtained using the method proposed in Section IV.2.
 The results of the five searches are aggregated and presented by the system.
@@
 These structures are not found in all the five CSP searches, while random search was able to find Cr2Ti twice and Mn2Ti once in five searches in the potential energy landscape of PFP.
-IV.2 CSP for binary and ternary systems
+IV.3 CSP for binary and ternary systems
 We demonstrated the capability of our CSP method using PFP to discover new crystal structure candidates for binary and ternary systems.
-The search conditions were the same as those in Section IV.1; however, the structures from MP are added to the initial population of the search.
+The search conditions were the same as those in Section IV.2; however, the structures from MP are added to the initial population of the search.
 The crystal structures obtained from the search were re-evaluated using DFT calculations to verify their validity.
 We report the results for systems In–Li, As–V, Al–Li–Pd, and La–Mo–O shown in Figs.
-6 and 7.
+8 and 9.
 Figure 8 depicts the convex hulls by our CSP methods and MP, respectively.
@@
 Crystal structures that update the convex hull of MP in the systems shown in Figs.
-6 and 7 are listed in Table 4 222 These crystal structures in xyz formats are available at Supplemental Materials.
+8 and 9 are listed in Table 4 222 These crystal structures in xyz formats are available at Supplemental Materials.
 .
@@
 Some of the listed crystal structures are visualized in Fig.
-8.
+10.
 Figure 8: Binary phase diagrams of (a) In–Li and (b) As–V systems.